40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 22 07:01:58 2022, Jc, Update, Coil Drivers, Adding Resistors and Reinstalling 8x
    Reply  Fri Apr 22 12:09:51 2022, Anchal, Update, Coil Drivers, Please update DCC pages 
    Reply  Wed Sep 7 14:32:15 2022, Anchal, Update, SUS, Updated SD coil gain to keep all coil actuation signals digitally same magnitude 
Message ID: 17134     Entry time: Wed Sep 7 14:32:15 2022     In reply to: 16802
Author: Anchal 
Type: Update 
Category: SUS 
Subject: Updated SD coil gain to keep all coil actuation signals digitally same magnitude 

The coil driver actuation strength for face coils was increased by 13 times in LO1, LO2, SR2, AS1, AS4, PR2, and PR3.

After the change the damping strenghth of POS, PIT, and YAW were reduced, but not SIDE coil output filter module requires higher digital input to cause same force at the output. This wouldn't be an issue until we try to correct for cross coupling at output matrix where we would want to give similar order of magnitude coefficients for SIDE coil as well. So I did the following changes to make input to all coil output filters (Face and side) same for these particular suspensions:

  • C1:SUS-LO1_SUSSIDE_GAIN 40.0 -> 3.077
    C1:SUS-AS1_SUSSIDE_GAIN 85.0 -> 6.538
    C1:SUS-AS4_SUSSIDE_GAIN 41.0 -> 3.154
    C1:SUS-PR2_SUSSIDE_GAIN 150.0 -> 11.538
    C1:SUS-PR3_SUSSIDE_GAIN 100.0 -> 7.692
    C1:SUS-LO2_SUSSIDE_GAIN 50.0 -> 3.846
    C1:SUS-SR2_SUSSIDE_GAIN 140.0 -> 10.769
  • C1:SUS-LO1_SDCOIL_GAIN -1.0 -> -13.0
    C1:SUS-AS1_SDCOIL_GAIN 1.0 -> 13.0
    C1:SUS-AS4_SDCOIL_GAIN -1.0 -> -13.0
    C1:SUS-PR2_SDCOIL_GAIN -1.0 -> -13.0
    C1:SUS-PR3_SDCOIL_GAIN -1.0 -> -13.0
    C1:SUS-LO2_SDCOIL_GAIN -1.0 -> -13.0
    C1:SUS-SR2_SDCOIL_GAIN -1.0 -> -13.0

In short, now 10 cts of input to C1:SUS-AS1_ULCOIL would create same actuation on UL as 10 cts of input to C1:SUS-AS1_SDCOIL will on SD.


In reply to

Quote: http://nodus.ligo.caltech.edu:8080/40m/16802

[Koji, JC]

Coil Drivers LO2, SR2, AS4, and AS1 have been updated a reinstalled into the system. 

LO2 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3        Unit: S2100008

LO2 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3                    Unit: S2100530

SR2 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3        Unit: S2100614

SR2 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3                    Unit: S2100615

AS1 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3        Unit: S2100610

AS1 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3                    Unit: S2100611

AS4 Coil Driver 1 (UL/LL/UR)now has R=100 // 1.2k ~ 92Ohm for CH1/2/3        Unit: S2100612

AS4 Coil Driver 2 (LR/SD)now has R=100 // 1.2k ~ 92Ohm for CH3                    Unit: S2100613

 

Quote: http://nodus.ligo.caltech.edu:8080/40m/16791

[JC Koji]

To give more alignment ranges for the SUS alignment, we started updating the output resistors of the BHD SUS coil drivers.
As Paco has already started working on LO1 alignment, we urgently updated the output Rs for LO1 coil drivers.
LO1 Coil Driver 1 now has R=100 // 1.2k ~ 92Ohm for CH1/2/3, and LO1 Coil Driver 2 has the same mod only for CH3. JC has taken the photos and will upload/update an elog/DCC.

We are still working on the update for the SR2, LO2, AS1, and AS4 coil drivers. They are spread over the workbench right now. Please leave them as they're for a while.
JC is going to continue to work on them tomorrow, and then we'll bring them back to the rack.

 

Quote: https://nodus.ligo.caltech.edu:8081/40m/16760

[Yuta Koji]

We took out the two coil driver units for PR3 and the incorrect arrangement of the output Rs were corrected. The boxes were returned to the rack.

In order to recover the alignment of the PR3 mirror, C1:SUS_PR3_SUSPOS_INMON / C1:SUS_PR3_SUSPIT_INMON / C1:SUS_PR3_SUSYAW_INMON were monitored. The previous values for them were {31150 / -31000 / -12800}. By moving the alignment sliders, the PIT and YAW values were adjusted to be {-31100 / -12700}. while this change made the POS value to be 52340.

The original gains for PR3 Pos/Pit/Yaw were {1, 0.52, 0.2}. They were supposed to be reduced to  {0.077, 0.04, 0.015}.
I ended up having the gains to be {0.15, 0.1, 0.05}. The side gain was also increased to 50.

----

Overall, the output R configuration for PR2/PR3 are summarized as follows. I'll update the DCC.

PR2 Coil Driver 1 (UL/LL/UR) / S2100616 / PCB S2100520 / R_OUT = (1.2K // 100) for CH1/2/3

PR2 Coil Driver 2 (LR/SD) / S2100617 / PCB S2100519 / R_OUT = (1.2K // 100) for CH3

PR3 Coil Driver 1 (UL/LL/UR) / S2100619 / PCB S2100516 / R_OUT = (1.2K // 100) for CH1/2/3

PR3 Coil Driver 2 (LR/SD) / S2100618 / PCB S2100518 / R_OUT = (1.2K // 100) for CH3

 

ELOG V3.1.3-